A uniform beam 3m long weighs 100N. Loads of 50 N and 150 N are placed on the beam at points which are ½ m and 2½ m, respectively, from its left end. If the beam is kept in a horizontal position by supports at its two ends, find these reactions. 


1
Expert's answer
2021-07-09T08:32:13-0400
"3R_2=1.5(100)+0.5(50)+2.5(150)\\R_2=183.3\ N"

"3R_1=1.5(100)+2.5(50)+0.5(150)\\R_1=350\ N"